OpenStudy (anonymous):

Determine f^-1(x) for f(x)= 3 times the square root of x-3.

OpenStudy (psymon):

\[3\sqrt{x-3}\] correct?

OpenStudy (anonymous):

yes! sorry, I didn't know how to do that.

OpenStudy (psymon):

You can click the equation button below the chat box thingy xD But alright, so if you need f inverse, then you need to switch x and y in your equation and then resolve for y. So basically you have: \[y = 3\sqrt{x-3}\] which becomes \[x = 3\sqrt{y-3}\] Do you know how to solve for y from here?

OpenStudy (anonymous):

I was pretty sure I did. I thought that since it's a cube root, you raise the whole function to the power of 3 to get rid of the cube root so it'd be x=(y-3)^3 right?